# Exploring Multiple Stellar Populations in Globular Clusters with Euclid: A Theoretical Overview and Insights from NGC 6397

### v1

@@
11institutetext: 1 Dipartimento di Fisica e Astronomia “Galileo Galilei”, Univ. di Padova, Vicolo dell’Osservatorio 3, Padova, IT-35122
2 Istituto Nazionale di Astrofisica - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova, IT-35122
@@ 1 Introduction
NGC 6397, the second-nearest GC, has been extensively studied in the context of multiple stellar populations. Observations have revealed two distinct stellar sequences along both the red-giant branch (RGB) and the MS, as identified in CMDs constructed using photometry from HST and JWST (Milone et al. 2012b, 2017b; Scalco et al. 2024b). These sequences correspond to chemically distinct populations, exhibiting moderate differences in helium, oxygen, and sodium abundances (Lind et al. 2011; Carretta et al. 2009; Milone et al. 2018; Marino et al. 2019a). The small elemental abundance differences between 1P and 2P stars, combined with the low metallicity of NGC 6397 ([Fe/H] $\sim-2.02$, based on the 2010 edition of the Harris (1996) catalog), result in only minor flux variations among stars with similar atmospheric parameters. This often makes it challenging to photometrically distinguish between 1P and 2P stars (e.g., Nardiello et al. 2015; Mehta et al. 2025).
@@
Date N $\times$ exposure time Mission Instrument Filter Program PI
Central Field RA=17h40m42s ${}^{{}^{\prime}}$58"
2006 May 29 1s$+$4$\times$15s HST WFC/ACS F606W 10775 A. Sarajedini
2006 May 29 1s$+$4$\times$15s HST WFC/ACS F814W 10775 A. Sarajedini
2014 March 27 1504s HST UVIS/WFC3 F275W 13297 G. Piotto
2014 June 11 1418s HST UVIS/WFC3 F275W 13297 G. Piotto
2014 March 27 620s HST UVIS/WFC3 F336W 13297 G. Piotto
2014 June 11 620s HST UVIS/WFC3 F336W 13297 G. Piotto
2014 March 27 66s HST UVIS/WFC3 F438W 13297 G. Piotto
2014 June 11 66s HST UVIS/WFC3 F438W 13297 G. Piotto
2016 March 03 4$\times$249s HST IR/WFC3 F110W 14124 M. Correnti
2016 March 03 4$\times$399s HST IR/WFC3 F160W 14124 M. Correnti
Field A RA=17h41m03s ${}^{{}^{\prime}}$38"
- April 08 126$\times$616-769s
- April 08 252$\times$584-804s
6$\times$1280s$+$12$\times$1405s
2014 March 27 683s HST WFC/ACS F475W 13297 G. Piotto
2014 March 27 610s HST WFC/ACS F814W 13297 G. Piotto
2023 March 14 12$\times$623s JWST NIRCam F150W2 1979 L. Bedin
2023 March 14 12$\times$623s JWST NIRCam F322W2 1979 L. Bedin
Field B RA=17h41m28s ${}^{{}^{\prime}}$10"
Field C RA=17h40m45s ${}^{{}^{\prime}}$32"
2003 May 16 327s$+$8$\times$400s HST WFC/ACS F775W 9480 J.,D. Rhodes
Field D RA=17h40m14s ${}^{{}^{\prime}}$37"
2014 June 11 640s HST WFC/ACS F475W 13297 G. Piotto
2014 June 11 567s HST WFC/ACS F814W 13297 G. Piotto
Field E RA=17h39m23s ${}^{{}^{\prime}}$26"
@@ 2.5 Results
The right panel of Figure 4 illustrates the $m_{\rm F814W}$ vs. $m_{\rm F814W}-m_{\rm F322W2}$ CMD for proper-motion-selected faint stars of NGC,6397 in Field A. The stellar sequence with colors exceeding $\sim 1.4$ represents the faintest part of the MS, extending down to the hydrogen-burning limit at $m_{\rm F814W}\sim 27$ mag, and includes the brightest brown dwarfs visible $m_{\rm F814W}\gtrsim 27$ mag. On the left, the white dwarf cooling sequence is prominently visible, showcasing notable slope changes around $m_{\rm F814W}\sim 26.5$ and 27.5 mag. These CMDs, along with the one shown in Figure 3, will be instrumental in identifying multiple populations among M-dwarfs and analyzing their internal proper motions. For earlier studies of the faintest stars in NGC 6397 based on similar data, we refer readers to Richer et al. (2006), Bedin et al. (2024), Gerasimov et al. (2024), and references therein.
@@ 4 Multiple populations in NGC 6397 from Euclid, HST and JWST
To investigate multiple stellar populations in NGC 6397 with Euclid data, we started analyzing the $H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ (left) and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$ CMDs plotted in Figure 9. In both diagrams, the color spread of stars fainter than the MS knee significantly exceeds the spread expected from observational errors, as evident from the comparison between the observed CMDs and the error bars. This excess color spread is consistent with the presence of multiple stellar populations characterized by different oxygen abundances.
@@
The left panel of Figure 10 presents the $\Delta_{\rm IE-JE}$ vs. $\Delta_{\rm YE-HE}$ ChM of M dwarfs within the magnitude range $21.019.8<H_{\rm E}<$ 21.0, where the impact of multiple populations on the MS color is most pronounced. This ChM, constructed following the methodology outlined by (Milone et al. 2017b, see also Section3), reveals two distinct stellar groups. Stars clustered near the origin of the reference frame correspond to the 1P, while the group located around ($\Delta_{\rm IE-JE}$, $\Delta_{\rm YE-HE}$) $\sim(-0.06,0.06)$ represents the 2P.
@@
Although the black isochrones provide a reasonable match to the observed CMDs, small discrepancies between the fiducial points of 1P stars and the black isochrone are evident. To mitigate the impact of these discrepancies on the estimation of the relative abundances of multiple populations in NGC,6397, we plotted in the insets of Figure 13 the $Y_{\rm E}-H_{\rm E}$ and $I_{\rm E}-J_{\rm E}$ color differences between the 2P and 1P fiducials at various $H_{\rm E}$ magnitudes. The blue and magenta lines in the insets represent the corresponding color differences between the O-poor isochrones and the black isochrone, showing a good match with the observed points. These isochrones are also superimposed on the ChM shown in the right panel of Figure 13, where they cluster into three distinct small regions marked with triangles. Our analysis indicates that isochrones with [O/Fe]=0.1 provide the best match for the bulk of 2P stars, while isochrones with [O/Fe]=0.0 likely represent the 2P stars occupying the extreme positions in the ChM.
@@ 4.1.1 Proper motion anisotropy
To study the internal dynamics of multiple populations in NGC 6397, we followed the procedure outlined in various works on this topic (e.g. Richer et al. 2013; Cordoni et al. 2020b, a; Libralato et al. 2022). First, we converted the proper motions from Cartesian coordinates ($\mu_{\mathrm{\alpha}}\cos\delta$, $\mu_{\mathrm{\delta}}$) to radial and tangential components ($\mu_{\mathrm{R}}$, $\mu_{\mathrm{T}}$). Here, radial proper motions are defined as positive when moving outward, while tangential proper motions are positive for counter-clockwise rotation.
@@ 5 Summary and conclusions
  • Both the $H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$ CMDs of proper-motion selected cluster members exhibit significant color broadening among M-dwarfs, which we attribute to the presence of multiple stellar populations. The $\Delta_{\rm YE-HE}$ vs. $\Delta_{\rm IE-JE}$ ChM constructed from these CMDs supports the existence of two stellar populations. Approximately 30% of the studied M-dwarfs belong to the 1P population, with pseudo-color broadening consistent with observational errors. In contrast, the 2P stars display an intrinsic color broadening, indicative of small but significant star-to-star chemical inhomogeneities.
@@
  • We compared the distribution of 1P and 2P stars in the Euclid CMDs ($H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$) and in the $\Delta_{\rm YE}-\Delta_{\rm HE}$ vs. $\Delta_{\rm IE-\rm JE}$ ChM with $\alpha$-enhanced isochrones from the BaSTi database (Pietrinferni et al. 2021), which account for the chemical composition of multiple populations in NGC,6397. The observations indicate that the bulk of 2P stars are depleted by $\sim$0.3 dex in [O/Fe] compared to the 1P stars, which exhibit [O/Fe]=0.4. Additionally, a small fraction of 2P stars, comprising about one-fifth of the total 2P population, show lower oxygen abundances with [O/Fe]$\sim$0.0.
@@
  • The wide field of view of Euclid has enabled an extensive investigation into the radial distribution of multiple populations in NGC,6397. Using Euclid photometry, we identified multiple populations among M-dwarfs spanning a radial range from $\sim$2 to $\sim$28 arcmin from the cluster center. In the crowded central regions, where disentangling 1P and 2P stars with Euclid data is challenging, we utilized the ChM of bright MS stars constructed from F275W, F336W, F438W, and F814W images obtained with the UVIS/WFC3 and WFC/ACS cameras on HST. Our analysis reveals that the fraction of 1P stars remains nearly constant across the studied radial interval. However, there is a subtle indication of a possible gradient, with the fraction of 2P stars increasing by less than $\sim$4when moving from the outermost cluster regions toward the center.
@@
Appendix. Multiple populations and cluster mass
As a byproduct of our research, we utilize the $m_{\rm F160W}$ vs.,$m_{\rm F110W}-m_{\rm F160W}$ CMD derived for central-field stars in NGC,6397, shown in the upper panel of Figure,19, to further investigate the relations identified by Dondoglio and collaborators. The results by Dondoglio et al. (2022) are reproduced in the bottom panels of Figure,19, where white and black dots represent the nine GCs studied by these authors and the open cluster NGC,6791, respectively.
The MS width of NGC,6397, determined from the CMD in Figure,19 using the method outlined by Dondoglio et al. (2022), is represented by an azure dot. The fact that our findings for NGC,6397 align with the established relations between cluster mass and oxygen variations supports the conclusion of Dondoglio and collaborators that the extent of internal oxygen variation in GCs depends on cluster mass. Similar trends involving internal nitrogen and helium variations among RGB stars have also been reported in previous studies (see Milone et al. 2017a, 2018; Lagioia et al. 2019; Jang et al. 2021).
acknowledgments

### v2

@@
11institutetext: 1 Dipartimento di Fisica e Astronomia “Galileo Galilei”, Univ. di Padova, Vicolo dell’Osservatorio 3, Padova, 35122, Italy
2 Istituto Nazionale di Astrofisica - Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, Padova, 35122, Italy
@@ 1 Introduction
NGC 6397, the second-nearest GC, has been extensively studied in the context of multiple stellar populations. Observations have revealed two distinct stellar sequences along both the red-giant branch (RGB) and the MS, as identified in CMDs constructed using photometry from HST and JWST (Milone et al. 2012b, 2017b; Scalco et al. 2024b). These sequences correspond to chemically distinct populations, exhibiting moderate differences in helium, oxygen, and sodium abundances (Lind et al. 2011; Carretta et al. 2009; Milone et al. 2018; Marino et al. 2019a). The small elemental abundance differences between 1P and 2P stars, combined with the low metallicity of NGC 6397 ([Fe/H] $\sim-2.02$, based on the 2010 edition of the Harris (1996) catalog), result in only minor flux variations among stars with similar atmospheric parameters. This often makes it challenging to photometrically distinguish between 1P and 2P stars (e.g., Nardiello et al. 2015; Branco et al. 2024; Mehta et al. 2025).
@@ 2.5 Results
The right panel of Figure 4 illustrates the $m_{\rm F814W}$ vs. $m_{\rm F814W}-m_{\rm F322W2}$ CMD for proper-motion-selected faint stars of NGC 6397 in Field A. The stellar sequence with colors exceeding $\sim 1.4$ represents the faintest part of the MS, extending down to the hydrogen-burning limit at $m_{\rm F814W}\sim 27$ mag, and includes the brightest brown dwarfs visible $m_{\rm F814W}\gtrsim 27$ mag. On the left, the white dwarf cooling sequence is prominently visible, showcasing notable slope changes around $m_{\rm F814W}\sim 26.5$ and 27.5 mag. These CMDs, along with the one shown in Figure 3, will be instrumental in identifying multiple populations among M-dwarfs and analyzing their internal proper motions. For earlier studies of the faintest stars in NGC 6397 based on similar data, we refer readers to Richer et al. (2006), Bedin et al. (2024), Gerasimov et al. (2024), and references therein.
@@ 4 Multiple populations in NGC 6397 from Euclid, HST and JWST
To investigate multiple stellar populations in NGC 6397 with Euclid data, we started analyzing the $H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ (left) and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$ CMDs plotted in Figure 9.
@@
The left panel of Figure 10 presents the $\Delta_{\rm IE-JE}$ vs. $\Delta_{\rm YE-HE}$ ChM of M dwarfs derived from these CMDs within the magnitude range $21.019.8<H_{\rm E}<$ 21.0, where the impact of multiple populations on the MS color is most pronounced. This ChM, constructed following the methodology outlined by (Milone et al. 2017b, see also Section3), reveals two distinct stellar groups. Stars clustered near the origin of the reference frame correspond to the 1P, while the group located around ($\Delta_{\rm IE-JE}$, $\Delta_{\rm YE-HE}$) $\sim(-0.06,0.06)$ represents the 2P.
@@
Although the black isochrones provide a reasonable match to the observed CMDs, small discrepancies between the fiducial points of 1P stars and the black isochrone are evident. To mitigate the impact of these discrepancies on the estimation of the relative abundances of multiple populations in NGC 6397, we plotted in the insets of Figure 13 the $Y_{\rm E}-H_{\rm E}$ and $I_{\rm E}-J_{\rm E}$ color differences between the 2P and 1P fiducials at various $H_{\rm E}$ magnitudes. The blue and magenta lines in the insets represent the corresponding color differences between the O-poor isochrones and the black isochrone, showing a good match with the observed points. These isochrones are also superimposed on the ChM shown in the right panel of Figure 13, where they cluster into three distinct small regions marked with triangles. Our analysis indicates that isochrones with [O/Fe]=0.1 provide the best match for the bulk of 2P stars, while isochrones with [O/Fe]=0.0 likely represent the 2P stars occupying the extreme positions in the ChM.
@@ 4.1.1 Proper motion anisotropy
To study the internal dynamics of multiple populations in NGC 6397, we followed the procedure outlined in various works on this topic (e.g. Richer et al. 2013; Cordoni et al. 2020b, a; Libralato et al. 2022; Ziliotto et al. 2025). First, we converted the proper motions from Cartesian coordinates ($\mu_{\mathrm{\alpha}}\cos\delta$, $\mu_{\mathrm{\delta}}$) to radial and tangential components ($\mu_{\mathrm{R}}$, $\mu_{\mathrm{T}}$). Here, radial proper motions are defined as positive when moving outward, while tangential proper motions are positive for counter-clockwise rotation.
@@ 5 Summary and conclusions
  • The $\Delta_{\rm YE-HE}$ vs. $\Delta_{\rm IE-JE}$ ChM constructed from the $H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$ CMDs of proper-motion selected cluster members exhibit significant color broadening among M-dwarfs, which is consistent with the existence of two stellar populations. Approximately 30% of the studied M-dwarfs belong to the 1P population, with pseudo-color broadening consistent with observational errors. In contrast, the 2P stars display an intrinsic color broadening, indicative of small but significant star-to-star chemical inhomogeneities.
@@
  • We compared the distribution of 1P and 2P stars in the Euclid CMDs ($H_{\rm E}$ vs. $I_{\rm E}-J_{\rm E}$ and $H_{\rm E}$ vs. $Y_{\rm E}-H_{\rm E}$) and in the $\Delta_{\rm YE}-\Delta_{\rm HE}$ vs. $\Delta_{\rm IE-\rm JE}$ ChM with $\alpha$-enhanced isochrones from the BaSTi database (Pietrinferni et al. 2021), which account for the chemical composition of multiple populations in NGC 6397. The observations indicate that the bulk of 2P stars are depleted by $\sim$0.3 dex in [O/Fe] compared to the 1P stars, which exhibit [O/Fe]=0.4. Additionally, a small fraction of 2P stars, comprising about one-fifth of the total 2P population, show lower oxygen abundances with [O/Fe]$\sim$0.0.
@@
  • The wide field of view of Euclid has enabled an extensive investigation into the radial distribution of multiple populations in NGC 6397. Using Euclid photometry, we identified multiple populations among M-dwarfs spanning a radial range from $\sim$2 to $\sim$28 arcmin from the cluster center. In the crowded central regions, where disentangling 1P and 2P stars with Euclid data is challenging, we utilized the ChM of bright MS stars constructed from F275W, F336W, F438W, and F814W images obtained with the UVIS/WFC3 and WFC/ACS cameras on HST. Our analysis reveals that the fraction of 1P stars remains nearly constant across the studied radial interval. However, there is a subtle indication of a possible gradient, with the fraction of 2P stars increasing by less than $\sim$4when moving from the outermost cluster regions toward the center.
@@
Acknowledgements.
@@
Appendix A HST and JWST images
The Table 3 provides the main information on the images of NGC 6397 collected by HST and JWST and used in the paper. In addition we provide the coordinates of each studied field.
Date N $\times$ exposure time Mission Instrument Filter Program PI
Central Field RA=17h40m42s ${}^{{}^{\prime}}$58"
2006 May 29 1s$+$4$\times$15s HST WFC/ACS F606W 10775 A. Sarajedini
2006 May 29 1s$+$4$\times$15s HST WFC/ACS F814W 10775 A. Sarajedini
2014 March 27 1504s HST UVIS/WFC3 F275W 13297 G. Piotto
2014 June 11 1418s HST UVIS/WFC3 F275W 13297 G. Piotto
2014 March 27 620s HST UVIS/WFC3 F336W 13297 G. Piotto
2014 June 11 620s HST UVIS/WFC3 F336W 13297 G. Piotto
2014 March 27 66s HST UVIS/WFC3 F438W 13297 G. Piotto
2014 June 11 66s HST UVIS/WFC3 F438W 13297 G. Piotto
2016 March 03 4$\times$249s HST IR/WFC3 F110W 14124 M. Correnti
2016 March 03 4$\times$399s HST IR/WFC3 F160W 14124 M. Correnti
Field A RA=17h41m03s ${}^{{}^{\prime}}$38"
- April 08 126$\times$616-769s
- April 08 252$\times$584-804s
6$\times$1280s$+$12$\times$1405s
2014 March 27 683s HST WFC/ACS F475W 13297 G. Piotto
2014 March 27 610s HST WFC/ACS F814W 13297 G. Piotto
2023 March 14 12$\times$623s JWST NIRCam F150W2 1979 L. Bedin
2023 March 14 12$\times$623s JWST NIRCam F322W2 1979 L. Bedin
Field B RA=17h41m28s ${}^{{}^{\prime}}$10"
Field C RA=17h40m45s ${}^{{}^{\prime}}$32"
2003 May 16 327s$+$8$\times$400s HST WFC/ACS F775W 9480 J.,D. Rhodes
Field D RA=17h40m14s ${}^{{}^{\prime}}$37"
2014 June 11 640s HST WFC/ACS F475W 13297 G. Piotto
2014 June 11 567s HST WFC/ACS F814W 13297 G. Piotto
Field E RA=17h39m23s ${}^{{}^{\prime}}$26"
Appendix B Multiple populations and cluster mass
As a byproduct of our research, we utilize the $m_{\rm F160W}$ vs.,$m_{\rm F110W}-m_{\rm F160W}$ CMD derived for central-field stars in NGC 6397, shown in the upper panel of Figure,19, to further investigate the relations identified by Dondoglio and collaborators. The results by Dondoglio et al. (2022) are reproduced in the bottom panels of Figure,19, where white and black dots represent the nine GCs studied by these authors and the open cluster NGC 6791, respectively.
The MS width of NGC 6397, determined from the CMD in Figure,19 using the method outlined by Dondoglio et al. (2022), is represented by an azure dot. The fact that our findings for NGC 6397 align with the established relations between cluster mass and oxygen variations supports the conclusion of Dondoglio and collaborators that the extent of internal oxygen variation in GCs depends on cluster mass. Similar trends involving internal nitrogen and helium variations among RGB stars have also been reported in previous studies (see Milone et al. 2017a, 2018; Lagioia et al. 2019; Jang et al. 2021).